I am having trouble with this question: "What is the product of (y+2)(y-8)? The answers available are
y^2+6y-16 , y^2-6y-16 , y^2-6y+16 , y^2+6y+16 . Please explain how you got the correct answer, thank you so much in advance , Emily.

Found 3 solutions by josgarithmetic, Theo, MathTherapy:
Answer by josgarithmetic(39617)

Perform the multiplication either using a lattice arrangement or using Distributive Property, and simplify.

Answer by Theo(13342)

all you need to do is multiply the factors together.
you would use the distributive law of multiplication.
that law says:

a * (b + c) =
(a*b) + (a*c)

that law also says:

(a + b) * (c + d) =
(a * (c + d)) + (b * (c + d)) =
(a*c) + (a*d) + (b*c) + (b*d).

in other words, every term in the multiplier needs to multiply every term in the multiplicand exactly one time.

this works for any number of terms.

a more complex example would be:

(a + b + c) * (d + e + f) =
(a * (d + e + f)) + (b * (d + e + f)) + (c * (d + e + f)) =
(a*d) + (a*e) + (a*f) + (b*d) + (b*e) + (b*f) + (c*d) + (c*e) + (c*f)

in your problem:

start with (y+2) * (y-8)
this becomes (y*(y-8)) + (2*(y-8))
which becomes (y*y) - (y*8) + (2*y) - (2*8)
simplify this to get y^2 - 8y + 2y - 16
combine like terms to get y^2 - 6y - 16

i believe that is one of your selections.






Answer by MathTherapy(10551)


I am having trouble with this question: "What is the product of (y+2)(y-8)? The answers available are
y^2+6y-16 , y^2-6y-16 , y^2-6y+16 , y^2+6y+16 . Please explain how you got the correct answer, thank you so much in advance , Emily.
Apply the FOIL method of distribution
(y + 2)(y - 8)
y(y - 8) + 2(y - 8)